$SPAD/src/input cardinal.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool cardinal.output
)set message test on
)set message auto off
)clear all
--S 1 of 16
(c0, c1, c2, c3, A0, A1): CardinalNumber
--R
--R                                                                   Type: Void
--E 1
--S 2 of 16
c0 := 0::NNI
--R
--R
--R   (2)  0
--R                                                         Type: CardinalNumber
--E 2
--S 3 of 16
c1 := 1::NNI
--R
--R
--R   (3)  1
--R                                                         Type: CardinalNumber
--E 3
--S 4 of 16
c2 := 2::NNI
--R
--R
--R   (4)  2
--R                                                         Type: CardinalNumber
--E 4
--S 5 of 16
c3 := 3::NNI
--R
--R

--R   (5)  3
--R                                                         Type: CardinalNumber
--E 5
--S 6 of 16
A0 := Aleph 0
--R
--R
--R   (6)  Aleph(0)
--R                                                         Type: CardinalNumber
--E 6
--S 7 of 16
A1 := Aleph 1
--R
--R
--R   (7)  Aleph(1)
--R                                                         Type: CardinalNumber
--E 7
--S 8 of 16
[finite? c2,    finite? A0]
--R
--R
--R   (8)  [true,false]
--R                                                          Type: List(Boolean)
--E 8
--S 9 of 16
[finite?  c2,    finite?  A0]
--R
--R
--R   (9)  [true,false]
--R                                                          Type: List(Boolean)
--E 9
--S 10 of 16
[countable? c2, countable? A0, countable? A1]
--R
--R
--R   (10)  [true,true,false]
--R                                                          Type: List(Boolean)
--E 10
--S 11 of 16
[c2 + c2, c2 + A1]
--R
--R
--R   (11)  [4,Aleph(1)]
--R                                                   Type: List(CardinalNumber)

--E 11
--S 12 of 16
[c2 - c1, c2 - c2, c2 - c3, A1 - c2, A1 - A0, A1 - A1]
--R
--R
--R   (12)  [1,0,"failed",Aleph(1),Aleph(1),"failed"]
--R                                   Type: List(Union(CardinalNumber,"failed"))
--E 12
--S 13 of 16
[c0 * c2, c1 * c2, c2 * c2, c0 * A1, c1 * A1, c2 * A1, A0 * A1]
--R
--R
--R   (13)  [0,2,4,0,Aleph(1),Aleph(1),Aleph(1)]
--R                                                   Type: List(CardinalNumber)
--E 13
--S 14 of 16
[c2**c0, c2**c1, c2**c2, A1**c0, A1**c1, A1**c2]
--R
--R
--R   (14)  [1,2,4,1,Aleph(1),Aleph(1)]
--R                                                   Type: List(CardinalNumber)
--E 14
--S 15 of 16
generalizedContinuumHypothesisAssumed true
--R
--R
--R   (15)  true
--R                                                                Type: Boolean
--E 15
--S 16 of 16
[c0**A0, c1**A0, c2**A0, A0**A0, A0**A1, A1**A0, A1**A1]
--R
--R
--R   (16)  [0,1,Aleph(1),Aleph(1),Aleph(2),Aleph(1),Aleph(2)]
--R                                                   Type: List(CardinalNumber)
--E 16
)spool
)lisp (bye)

———-

References^{[1] nothing}

6